# The temperature and metallicity distributions of the ICM: insights with TNG-Cluster for XRISM-like observations

### v1

@@ Abstract
The new era of high-resolution X-ray spectroscopy will significantly improve our understanding of the intra-cluster medium (ICM) by providing precise constraints on its underlying physical properties. However, spectral fitting requires reasonable assumptions on the thermal and chemical distributions of the gas. We use the output of TNG-Cluster, the newest addition to the IllustrisTNG suite of cosmological magnetohydrodynamical simulations, to provide theoretical expectations for the multi-phase nature of the ICM across hundreds of clusters (M) based upon a realistic model for galaxy formation and evolution. We create and analyse, in an observer-like manner, end-to-end XRISM/Resolve mock observations towards cluster centres. We then systematically compare the intrinsic properties of the simulated gas with the inferred ones from spectral fitting via a variety of commonly used spectral-emission models. Our analysis suggests that models with a distribution of temperatures, such as bvlognorm and bvgadem, better describe the complex thermal structure of the ICM, as predicted by TNG-Cluster, but incur biases of keV (16th-84th percentiles). The 1T bvapec is too simplistic for the predicted broad temperature distributions, while a 2T double bvapec model systematically fails to capture the input temperature structure. However, all spectral emission models systematically underestimate the Fe abundance of the central ICM by solar ( 20 per cent) primarily due to projection effects. Selecting only strong cool core clusters leads to minor improvements on inference quality, removing the majority of outliers but maintaining similar overall biases and cluster-to-cluster scatter.
@@
Galaxy clusters are excellent probes of the large-scale structure of the Universe due to their rarity, masses and preferential position at the knots of the cosmic web (e.g. Allen et al., 2011). Their extended gaseous atmospheres consisting of X-ray emitting plasma, known as the intra-cluster medium (ICM), play a key role in the evolution of large structures, since processes such as mergers, gas accretion from the cosmic web and feedback all leave signatures in the distribution and properties of this gas.
Various X-ray telescopes (e.g. XMM-Newton, Chandra, Suzaku, SRG/eROSITA) have studied the ICM and its thermal and chemical properties across spatial and mass scales (e.g. Kravtsov & Borgani, 2012; Mernier et al., 2018; Walker et al., 2019; Mernier et al., 2023, and references therein). However, the low CCD spectral resolution (120 eV) of the detectors so far limits the scope of X-ray analyses. For example, at CCD resolution, the dynamical state of the gas is hard to assess, as direct measurements of the gas motions cannot be resolved but for very high speeds; similarly, neighbouring emission lines may get blended, complicating the recovery of e.g. the abundances of various elemental species and the extent of the multi-phase thermal structure of the gas. As a result, many simplifying assumptions, such as coupling individual abundances to Iron (Fe) or assuming single temperature gas, have been routinely made.
@@
However, by overcoming many previous degeneracies, systematic uncertainties become a more prominent source of errors. Assumptions made during the analysis of the X-ray spectra – such as the assumed shape of the temperature and metallicity distributions of the ICM or on the atomic libraries describing the emission lines – as well as unavoidable projection effects – all impact inferences, and may lead to significant biases. This can be problematic, especially for science cases that are sensitive to the exact values of cluster properties, such as their metal budgets (e.g. Gastaldello et al., 2021, and references therein) and the X-ray observable-mass relations used for cosmological inference (e.g. most recently Ghirardini et al., 2024).
@@
Unfortunately, the physical properties of the ICM are sufficiently complex that these issues cannot be assessed empirically. In this case, numerical simulations are useful tools. With the inclusion of ever more and more realistic physical processes, simulations provide an increasingly detailed view of the ICM (e.g. Nelson et al., 2024, and references therein). For example, in past works, simulations of clusters have quantified biases associated with observations, such as the differences between emission- vs. mass-weighted metallicity profiles or the effects of projection effects (e.g. Biffi et al., 2018; Veronesi et al., 2024). However, no analysis has yet quantified the impact of specific parametric model choices, of e.g. the ICM temperature distribution, when fitting X-ray spectra.
@@
In this work, we aim to assess the systematic effects associated with the inference of the ICM thermal properties and Fe abundance by using varying and complementary spectral-emission models that are routinely adopted in the field. To do so, we use TNG-Cluster (Nelson et al., 2024), a recent extension of the IllustrisTNG simulations,222www.tng-project.org which returns a rich and statistically-large sample of simulated galaxy clusters well suited to study the ICM. Based on the IllustrisTNG model (Weinberger et al., 2017; Pillepich et al., 2018a) applied to a suite of zoom-in cosmological magneto-hydrodynamical simulations, TNG-Cluster not only creates galaxies at that are consistent with observations, but also – and of relevance to this work – produces the ICM in clusters with overall characteristics that are consistent with observations (Nelson et al., 2024).
[…]
@@
This paper is arranged as follows. In Section 2.1 we present our sample of simulated galaxy clusters from TNG-Cluster. In Section 2.2 we describe our end-to-end mocking procedure, used to create XRISM/Resolve like observations of our input clusters, and to analyse them. In Section 3.1 we first examine the intrinsic properties of the ICM, as predicted by TNG-Cluster. We then compare the inferences of various spectral modelling approaches on the temperature distribution and Fe abundance in Sections 3.2 and 3.3, respectively. We discuss the reasons and implications of our findings and expand on various effects in Section 4, and summarize in Section 5.
@@ 2.1 The TNG-Cluster simulation suite
TNG-Cluster is a collection of 352 zoom-in magnetohydrodynamical simulations of galaxy clusters drawn from a 1 Gpc parent box, with halo masses of log i.e. log. It complements the volume-limited TNG300 simulation (Springel et al., 2018; Pillepich et al., 2018b; Nelson et al., 2018; Naiman et al., 2018; Marinacci et al., 2018) with the highest halo masses. In particular, a subset of halos with log10M at have been randomly chosen for re-simulation in TNG-Cluster in order to produce a uniform selection, and hence to compensate for the halo mass drop in TNG300, while simultaneously including all halos with log10(M. For a more detailed description of the halo selection we refer to Nelson et al. (2024). Here we summarize additional salient aspects.
TNG-Cluster was simulated with the AREPO code (Springel, 2010) at a resolution of 1.1 M⊙ in gas/stellar target mass, matching that of the TNG300 fiducial run. It is fully based on the IllustrisTNG model (Weinberger et al., 2017; Pillepich et al., 2018a). This includes a variety of processes that are needed to simulate galaxy formation and evolution, such as stellar formation and evolution, metal enrichment from AGB stars, type Ia and type II supernovae by tracking nine elements (H, He, C, N, O, Ne, Mg, Si, and Fe) in addition to Europium, radiative cooling/heating, seeding and growth of SMBH and both stellar and SMBH (thermal and kinetic) feedback.
@@ 2.2.1 Modeling the intrinsic X-ray signal
First, we model the X-ray emission of the ICM based on the thermodynamical and chemical properties of the simulation gas cells, as predicted by TNG-Cluster. For the purposes of this work, we use the 4.4.0 version of the PyXSIM package (ZuHone et al., 2014) in combination to the SOXS package (ZuHone et al., 2023).
Similar to Truong et al. (2024), for every cluster, we consider as emitting sources all the gas cells that are within the Friend-of-Friend halo and that are within a 3D radius of R. Throughout, the centre of a cluster from TNG-Cluster is defined as the minimum of the potential. Furthermore, we assume that all gas cells satisfying the following criteria contribute to the X-ray emission of a given targeted cluster: have a temperature higher than K ( 0.03 keV), a negative net cooling rate, and are not star forming.
@@
The properties of the gas cells predicted within the TNG-Cluster simulation and relevant for this pipeline are: density, temperature, and absolute elemental abundances (w.r.t. Hydrogen). More specifically, for the latter, we use the TNG-Cluster outcome for all nine elements that the simulation tracks and a relative abundance ratio of for the rest (adopting Asplund et al., 2009), assuming that the absolute Fe abundance is a reasonable proxy of the underlying metal enrichment and that solar abundance ratios hold for those elements. Additionally, we account for the 3D position and velocity of each gas cell also contribute to the properties of the generated X-ray photons. This allows us to account for effects such as thermal broadening and Doppler shift of the emission lines, producing an accurate representation of the X-ray emission from the ICM of our simulated galaxy clusters.
@@ 2.2.2 Observational strategy
Every cluster is observed with one pointing, focusing on the centre, and hence covering an area given by XRISM Field of View (FoV): 3’x3’. In order to best approximate realistic observations of our targets, we also apply XRISM/Resolve’s instrumental responses to the intrinsic incoming photons. Namely, we use the Cycle 1 response files, as they have been incorporated in SOXS, for a high resolution ( 5eV) response matrix file (RMF) and a standard point-source ancillary response file (ARF) with the gate valve closed.
@@
Throughout this work, we utilize four commonly-used spectral-emission models with different assumptions on the thermal structure of gas. The models in question are:
@@
double bvapec: this consists of two bvapec models and is primarily used to describe two distinct phases of gas (hot and cooler) assuming de facto a binomial thermal structure of the ICM;
@@
For the Fe abundance, all adopted models assume that the observed ICM is described by a single value, as is the case for temperature with the bvapec and double bvapec models.
We note that for all cases, both the line broadening due to the kinematics of the gas and the abundances of Si, S, Ar, Ca and Ni are also taken into consideration, are modelled independently, and are simultaneously fit for – however, further analysis and their exploitation are beyond the scope of this work.
Regarding our spectral fitting strategy, all our extracted spectra are fitted in the [1.8-10] keV range. Even though the target redshift is known, we allow for the fit to initially vary it over a 50 per cent range in order to account for possible bulk motions of the gas along our line of sight. We use a guided approach to fit the model parameters gradually. However, it is still possible for the model to produce a non-valid fit. For this reason we introduce convergence criteria.
2.3.1 Acceptance criteria for spectra best fits
@@
For both the bvgadem and bvlognorm models, we require that the uncertainties on the central temperature and on the width of the distributions are smaller than the best-fit values themselves, assuming a signal-to-noise ratio of at least 1 as a minimum requirement for convergence. For both models, we also require that the width of the assumed distribution is smaller than the best-fit central temperature, as otherwise the model would explicitly return negative temperatures to describe the thermal state of the gas. For the double bvapec model, we apply this same convergence criteria to each of the two temperature components, in order for the model to be meaningfully different from the single bvapec one.
If a best-fit result fails any of the above criteria, the fit is considered rejected; if it fulfils all of them, it is accepted. We assume by construction that the bvapec model is always accepted.
@@ 3 Results
Firstly, we compare the outputs of the spectral-emission models to the mass-weighted temperature and Fe abundance of the gas in the simulation located in a spherical volume centred at the centre of the cluster, with a radius of 1.5’ (33.5 kpc at the redshift of Perseus, to approximate XRISM FoV): we refer to this as gas in the central sphere or in the 3D cluster centres. Our choice is motivated by the fact that, observationally, one would hope that the output of the extracted spectrum represents the local conditions of the de-projected gas in the region of interest – for us, the XRISM FoV at the cluster centre.
We also compare the spectral model inferences to the emission-weighted properties of all the gas in the adopted FoV for a central pointing, namely also considering all the foreground and background ICM gas: we refer to this as gas in a central pointing with a XRISM-like FoV. Previous studies have shown that the observed ICM abundance profiles can be better reproduced by weighting the properties of gas particles with emission rather than mass (Biffi et al., 2017, 2018). Therefore, this can be considered a more apt comparison with our observables and a characterization of the “true” ICM state that is by construction closer to what the spectral fitting can return.
@@
where , the electron and hydrogen number densities and the volume of the gas cell, tracked by the simulation.

### v2

@@ Abstract
The new era of high-resolution X-ray spectroscopy will significantly improve our understanding of the intra-cluster medium (ICM) by providing precise constraints on its underlying physical properties. However, spectral fitting requires reasonable assumptions on the thermal and chemical distributions of the gas. We use the output of TNG-Cluster, the newest addition to the IllustrisTNG suite of cosmological magnetohydrodynamical simulations, to provide theoretical expectations for the multi-phase nature of the ICM across hundreds of clusters (M) based upon a realistic model for galaxy formation and evolution. We create and analyse, in an observer-like manner, end-to-end XRISM/Resolve mock observations towards cluster centres. We then systematically compare the intrinsic temperature and Fe abundance of the simulated gas with the inferred ones from spectral fitting via a variety of commonly used spectral-emission models. Our analysis suggests that models with a distribution of temperatures, better describe the broad thermal distributions of the ICM, as predicted by TNG-Cluster, but still incur biases in the inferred temperature of keV (16th-84th percentiles). However, all spectral-emission models systematically underestimate the Fe abundance of the central ICM by 0.12 Solar ( 22 per cent), almost an order of magnitude higher than the abundance errors reported in the literature, primarily due to projection effects. Selecting only strong cool core clusters leads to minor improvements on inference quality, removing the majority of outliers but maintaining similar overall biases and cluster-to-cluster scatter.
@@
Galaxy clusters are excellent probes of the large-scale structure of the Universe due to their rarity, masses and preferential position at the knots of the cosmic web (e.g. Allen et al., 2011). Their extended gaseous atmospheres consisting of X-ray emitting plasma, known as the intra-cluster medium (ICM), play a key role in our understanding of the evolution of large structures, since processes such as mergers, gas accretion from the cosmic web and feedback all leave signatures in the distribution and properties of this gas.
Various X-ray telescopes (e.g. XMM-Newton, Chandra, Suzaku, SRG/eROSITA) have studied the ICM and its thermal and chemical properties across spatial and mass scales (e.g. Kravtsov & Borgani, 2012; Mernier et al., 2018; Walker et al., 2019; Mernier et al., 2023, and references therein). However, the low spectral resolution of the CCD detectors (120 eV) so far limits the scope of X-ray analyses. For example, at CCD resolution, the dynamical state of the gas is hard to assess, as it lacks the spectral resolution for a direct measurement of gas motions from the Doppler shifts of spectral lines. Similarly, neighbouring emission lines may get blended, complicating the recovery of e.g. the abundances of various elemental species and the extent of the multi-phase thermal structure of the gas. As a result, many simplifying assumptions, such as coupling individual abundances to Iron (Fe) or assuming single temperature gas, have been routinely made.
@@
However, by overcoming many previous degeneracies, systematic uncertainties become a more prominent source of errors. Assumptions made during the analysis of the X-ray spectra – such as the assumed shape of the temperature and metallicity distributions of the ICM or on the atomic libraries describing the emission lines, as well as unavoidable projection effects – all impact inferences, and may lead to significant biases. This can be problematic, especially for science cases that are sensitive to the exact values of cluster properties, such as their metal budgets (e.g. Gastaldello et al., 2021, and references therein) and the X-ray observable-mass relations used for cosmological inference (e.g. most recently Ghirardini et al., 2024).
@@
Unfortunately, the physical properties of the ICM are sufficiently complex that these issues cannot be assessed empirically. In this case, numerical simulations are useful tools. With the inclusion of ever more and more realistic physical processes, simulations provide an increasingly detailed view of the ICM (e.g. Nelson et al., 2024, and references therein). A number of previous works have quantified biases on the observationally inferred properties of clusters by mocking the X-ray emission of simulated clusters and by focusing on past, current and future X-ray missions (e.g. Gardini et al., 2004; Rasia et al., 2006, 2008; Biffi et al., 2018; Cucchetti et al., 2018; Roncarelli et al., 2018; Mernier et al., 2020; Barnes et al., 2021; ZuHone et al., 2023b; Castellani et al., 2024; Veronesi et al., 2024; Zhang et al., 2024; Sanders, 2023, and references therein). However, no analysis has yet quantified the systematic effects due to specific parametric model choices, such as the ICM temperature distribution, when fitting high-resolution X-ray spectra.
@@
In this work, we aim to assess the ability of spectral-emission models that are routinely adopted in the field to infer the mass-weighted temperature and Fe abundance of the ICM in cluster centres. Specifically, we attempt to quantify the systematic errors that might be introduced in said inferences due to the assumed description of the ICM’s temperature distribution, by simultaneously accounting for cluster-to-cluster variations. To do so, we use TNG-Cluster (Nelson et al., 2024), a recent extension of the IllustrisTNG simulations,222www.tng-project.org which returns a rich and statistically-large sample of simulated galaxy clusters well suited to study the ICM. Based on the IllustrisTNG model (Weinberger et al., 2017; Pillepich et al., 2018a) applied to a suite of zoom-in cosmological magneto-hydrodynamical simulations, TNG-Cluster not only creates galaxies at that are consistent with observations, but also – and of relevance to this work – produces the ICM in clusters with overall characteristics that are consistent with observations (Nelson et al., 2024).
It is thanks to the combination of numerical resolution and the adopted physical modelling that the output of TNG-Cluster is reasonably, if not quantitatively, realistic. This includes a mix of strong, weak and non cool cores (Lehle et al., 2024, 2025), realistic levels of gas velocity dispersion (Truong et al., 2024) and kinematics (Ayromlou et al., 2024), a non-negligible amount of cool gas (Rohr et al., 2024; Staffehl et al., 2025), a diversity of radio relics (Lee et al., 2024), and rich X-ray morphologies similar to observations (Chadayammuri et al., 2024), including cold fronts, shocks, and cavities (Truong et al., 2024; Prunier et al., 2024). Additionally, the ICM returned by TNG-Cluster exhibits thermodynamical profiles across the 352 clusters – density, temperature, entropy, and pressure – that are in the ballpark of observational inferences (Lehle et al., 2024). The distribution of cooling radii, derived from spectral fitting on both TNG-Cluster mocks and real Chandra observations, are quantitatively consistent (Prunier et al., 2025). And the metallicity profiles appear compatible with available observational results taken at face value (Nelson et al., 2024). Overall, the realism of the ICM produced by TNG-Cluster further supports the credibility of the analysis we set out.
@@
This paper is arranged as follows. In Section 2.1 we present our sample of simulated galaxy clusters from TNG-Cluster. In Section 2.2 we describe our end-to-end mocking procedure, used to create XRISM/Resolve like observations of our input clusters, and to analyse them. In Section 3.1 we first examine the intrinsic properties of the ICM, as predicted by TNG-Cluster. We then compare the inferences of various spectral-emission modelling approaches on the temperature distribution and Fe abundance in Sections 3.2 and 3.3, respectively. We discuss the reasons and implications of our findings and expand on various effects in Section 4, and summarize in Section 5.
@@ 2.1 The TNG-Cluster simulation suite
TNG-Cluster is a collection of 352 zoom-in magnetohydrodynamical simulations of galaxy clusters with halo masses of log i.e. log. These are drawn from a cubic parent box of about 1 comoving Gpc per side. It complements the volume-limited TNG300 simulation (Springel et al., 2018; Pillepich et al., 2018b; Nelson et al., 2018; Naiman et al., 2018; Marinacci et al., 2018) with the highest halo masses. In particular, a subset of halos with log at have been randomly chosen for re-simulation in TNG-Cluster in order to produce a uniform selection, and hence to compensate for the halo mass drop in TNG300, while simultaneously including all halos with log from the same parent box. The volumes for re-simulations have been chosen such that the high-resolution zoom-in region around each cluster has a radius typically larger than two times the halo virial radius. For a more detailed description of the halo selection and model choices we refer to Nelson et al. (2024). Here we summarize additional salient aspects.
TNG-Cluster was simulated with the AREPO code (Springel, 2010) at a resolution of 1.1 M⊙ in gas/stellar target mass, matching that of the TNG300 fiducial run. It is fully based on the IllustrisTNG model (Weinberger et al., 2017; Pillepich et al., 2018a). This includes a variety of processes that are needed to simulate galaxy formation and evolution, such as stellar formation and evolution, metal enrichment, radiative cooling/heating, seeding and growth of SMBH and both stellar and SMBH (thermal and kinetic) feedback. Specifically, the IllustrisTNG model produces stellar particles, each representing a single-age stellar population with an initial mass function described by Chabrier (2003). Such populations evolve and return metals to the surrounding inter-stellar medium via AGB stars, type Ia and type II supernovae over time (Vogelsberger et al., 2013; Wiersma et al., 2009) based on the stellar yields described in detail in Pillepich et al. (2018a). The simulation directly tracks nine elements (H, He, C, N, O, Ne, Mg, Si, and Fe), in addition to Europium, tagging their production channel.
@@ 2.2.1 Modeling the intrinsic X-ray signal
First, we model the X-ray emission of the ICM based on the thermodynamical and chemical properties of the simulation gas cells, as predicted by TNG-Cluster. For the purposes of this work, we use the 4.4.0 version of the PyXSIM package (ZuHone et al., 2015), which follows the approach of the PHOX code (Biffi et al., 2012, 2013), in combination to the SOXS package (ZuHone et al., 2023a).
Similar to recent works (e.g. ZuHone et al., 2023b; Truong et al., 2024), for every cluster, we consider all the gas cells that are within a 3D radius of , including contribution from satellites, as emitting sources. While this approach excludes emission from structures outside that radius, ZuHone et al. (2023b) has shown for eROSITA mocks that their effect on the inferred cluster properties is negligible. Throughout, the centre of a cluster from TNG-Cluster is defined as the minimum of the potential. Furthermore, we assume that all gas cells satisfying the following criteria contribute to the X-ray emission of a given targeted cluster: have a temperature higher than K ( 0.03 keV), a negative net cooling rate, and are not star forming.
@@
The properties of the gas cells predicted within the TNG-Cluster simulation and relevant for this pipeline are: density, temperature, and absolute elemental abundances (w.r.t. Hydrogen). More specifically, for the latter, we use the TNG-Cluster outcome for all nine elements that the simulation tracks and a relative abundance ratio of for the rest (adopting Asplund et al., 2009), assuming that the absolute Fe abundance is a reasonable proxy of the underlying metal enrichment and that solar abundance ratios hold for those elements. Additionally, we account for the 3D position and velocity of each gas cell, to properly account for their effects to the properties of the generated X-ray photons. Namely, effects such as thermal broadening and Doppler shift of the emission lines are explicitly modelled by PyXSIM, producing an accurate representation of the X-ray emission from the ICM of our simulated galaxy clusters.
@@ 2.2.2 Observational strategy
Every cluster is observed with a single pointing aimed at each cluster’s centre and hence detecting the emission from all gas cells in a box volume: this hence represents a subset of the original ( 4-5 pMpc) re-simulation volume of the zoom simulations and is meant to replicate the XRISM/Resolve’s Field of View (FoV). In order to best approximate realistic observations of our targets, we also apply XRISM/Resolve’s instrumental responses to the intrinsic incoming photons. Namely, we use the Cycle 1 response files, as they have been incorporated in SOXS, for a high resolution ( 5eV) response matrix file (RMF) and a standard point-source ancillary response file (ARF) with the gate valve closed.
@@
Throughout this work, we utilize four commonly-used spectral-emission models with different assumptions on the thermal structure i.e. temperature distribution of gas. The models in question are:
@@
double bvapec: this consists of two bvapec models and is primarily used to describe two distinct phases of gas (hot and cooler) assuming de facto a bimodal temperature distribution of the ICM;
@@
For the Fe abundance, all adopted models assume that the observed ICM is described by a single value, as is the case for temperature with the bvapec and double bvapec models. We note that for all cases, both the line broadening due to the kinematics of the gas and the abundances of Si, S, Ar, Ca and Ni are taken into consideration, are modelled independently, and are simultaneously fit for. This is motivated by the fact that, if left fixed at a specific values, degeneracies between parameters could influence the model inferences on the temperature and Fe abundance. Consistently with Truong et al. (2024) and Xrism Collaboration et al. (2025c), our inferred velocity dispersions are comparable to the values obtained by central XRISM pointing observations (see Figure 15). However, further analysis and their exploitation are part of a future work. We note that H/He is kept fixed at 1 Solar and the abundances of all remaining elements are coupled to Fe, since they don’t have emission lines in our energy range.
For the purposes of this analysis, we do not deproject our spectra. With this choice we remain consistent with the spectral analysis of most recent XRISM observations (e.g. Xrism Collaboration et al., 2025a, b; Rose et al., 2025; Fujita et al., 2025). The only exception is that of XRISM Collaboration et al. (2025) for the Centaurus cluster, where a deprojected profile has been used for the interpretation of the results and has been constructed by fitting previous Chandra observations of Centaurus. An exact replication of this particular approach and study would require the full mocking and spectral fitting of Chandra-like observations for our entire sample, which would be beyond the scope of this project. Therefore, we choose to not account for projection effects on our spectral fitting, but we expand more on the implications of this choice in Section 4.2.1.
Regarding our spectral fitting strategy, all our extracted spectra are fitted in the [1.8-10] keV range. Even though the target redshift is known, we allow for the fit to vary it over a 50 per cent range in order to account for possible bulk motions of the gas along our line of sight. We use a guided approach to fit the model parameters gradually. However, it is still possible for the model to produce a non-valid fit. For this reason we introduce convergence criteria.
2.3.1 Acceptance criteria for spectral best fits
@@
For both the bvgadem and bvlognorm models, we require that the uncertainties on the central temperature and on the width of the distributions are smaller than the best-fit values themselves, as a minimum requirement for convergence. For the bvgadem model, we also require that the width of the assumed distribution is smaller than the best-fit central temperature, as otherwise the model would either imply the presence of negative temperatures or, due to its hard limit at 0, significantly alter its shape, in order to describe the thermal state of the gas. We note that this criterion is not as significant as the convergence of individual best-fit parameters, with 1 per cent of our clusters being rejected due to it. For the double bvapec model, we apply this same convergence criteria to each of the two temperature components, in order for the model to be meaningfully different from the single bvapec one. Lastly, we also apply a goodness-of-fit criterion to our spectra, requiring the C-statistic (Cash, 1979) of our spectral fits to deviate less than 3 from its expected Cash-statistic C value, following Kaastra (2017). If a best-fit result fails any of the above criteria, the fit is considered rejected; if it fulfils all of them, it is accepted.
@@ 3 Results
Firstly, we compare the output temperature and Fe abundance of the best-fit spectral-emission models to the mass-weighted temperature and Fe abundance of the gas in the simulation located in a spherical volume centred at the centre of the cluster, with a radius of 1.5’ (33.5 kpc at the redshift of Perseus, to approximate XRISM FoV): we refer to this as gas in the central sphere or in the 3D cluster centre. Our choice is motivated by the fact that, observationally, one would hope that the output of the extracted spectrum represents the local conditions of the de-projected gas in the region of interest – for us, the XRISM FoV at the cluster centre.
Secondly, we compare the spectral-emission model inferences to the emission-weighted properties of all the gas in the adopted FoV for a central pointing, namely also considering all the foreground and background ICM gas: we refer to this as gas in a central pointing with a XRISM-like FoV. Previous studies have shown that the observed ICM abundance profiles can be better reproduced by weighting the properties of gas particles/cells with emission rather than mass (Biffi et al., 2017, 2018). Therefore, this can be considered a more apt comparison with our observables and a characterization of the “true” ICM state that is by construction closer to what the spectral fitting can return.
@@
where , are the electron and hydrogen number densities and the volume of the gas cell, tracked by the simulation. We note that this weighting scheme differs from emission-weighted schemes that use the emissivity of each gas cell as their weight. Our choice in the weighting scheme is motivated by the fact that EM is indicative of each gas cell’s contribution to the spectrum, as the basis of the spectral-emission model’s normalisation. Furthermore, Cucchetti et al. (2018) have indicated that the metallicity of the gas, weighted by its EM, can be accurately recovered by fitting mock, high-resolution spectra.